# Quantum Simulation of Noisy Quantum Networks

### v1

@@ V Network simulation with noisy quantum devices
In any quantum computing hardware, NISQ devices in particular, noise and decoherence are inevitable factors that affect the reliability of quantum computations. When simulating quantum communication or network processes on such devices, it becomes imperative to incorporate these noisy characteristics. Achieving a faithful simulation of quantum communication processes on quantum hardware necessitates the utilization of diverse tools and techniques. These tools enable the replication of quantum schemes within a quantum circuit while effectively managing -or even leveraging- the inherent noise of the system. As already stressed before, our evaluations typically focus on simulating specific quantum channels, with a designated target denoted as ^$\hat{\mathcal{E}}_{\text{t}}$, representing the desired channel to be implemented on the quantum hardware. While an ideal, noiseless quantum computer could implement the channel with perfect fidelity using the techniques discussed in Sec. IV, imperfect quantum hardware necessitates the development of strategies to mitigate noise effects, exploit them, or transform certain quantum channels. We primarily rely on static noise models. In practice, however, quantum device noise often exhibits temporal fluctuations due to environmental factors such as temperature drift, vibrations, and electromagnetic interference. While our framework does not simulate noise dynamically at every intermediate time step, it remains applicable by focusing on reproducing the effective initial-to-final quantum process, which can be captured as a single CP map. In this context, we propose several general methods to address these challenges and evaluate their effectiveness through relevant examples.
@@
The first strategy we propose is denoted as the building-block method, illustrated in Fig. 4a. In this method, the circuits used to simulate quantum channels are treated as fixed building blocks. These circuits, representing the fundamental elements of the simulation process, are not directly modifiable, including the sequences of gates and operations. However, it is possible to introduce additional channels within the quantum circuit framework. These supplementary channels, also susceptible to noise, are strategically applied in various configurations. For instance, they can be applied before an input channel, after it, or both (referred to as interleaved), with the aim of achieving a satisfactory approximation of the desired target channel ^${\hat{\mathcal{E}}_{\text{t}}}$. Note the target channel equals the noiseless implementation of the input channel. Additionally, the method allows for the consideration of non-trivial classical mixtures of these channels applied with different probabilities. It is crucial to note that directly inverting a quantum channel is not possible.
@@
^$\hat{\mathcal{E}}_{\text{out}}=\sum_{i,j}p_{ij}\,\hat{\zeta}^{\prime}_{i}\circ\hat{\mathcal{E}}^{\prime}_{\text{t}}\circ\hat{\xi}^{\prime}_{j},$ (16)
such that maximizes the channel fidelity between the output and target channels $F(\hat{\mathcal{E}}_{\text{out}},\hat{\mathcal{E}}_{t})$, such that ^$\hat{\mathcal{E}}_{\text{out}}\approx\hat{\mathcal{E}}_{\text{t}}$, where $\{p_{ij}\}$ defines a probability distribution and $\{\hat{\zeta}_{i},\hat{\xi}_{j}\}$ are quantum channels. The prime notation, i.e. $\hat{\xi}^{\prime}$ indicates the channel implementation is subjected to computer noise that deviates from the ideal channel implementation $\hat{\xi}$. This deviation can be considered by the method when optimizing over the construction of the channels $\{\hat{\zeta}_{i},\hat{\xi}_{j}\}$. In the examples shown below we model the hardware noise using the models introduced in Sec. III.2.
@@
Instead of solely compensating for noise, this method can also be utilized for channel transformation, with the objective of converting some input channel ^$\hat{\mathcal{E}}{\text{in}}$ into a –in principle very different– target channel ^$\hat{\mathcal{E}}{\text{t}}$ using the aforementioned techniques. In the context of large network simulations, there may arise scenarios where a specific type of noise needs to be simulated from a particular point. This strategy enables the direct transformation of the noise present in the simulation into the desired form, effectively harnessing the inherent noise of the quantum computer. From a fundamental perspective, the task of transforming one type of noise into another using additional channels applied before and after holds significant importance, even within the context of noiseless quantum hardware assumptions.
@@
Fig. 5: (a) Channel infidelity $(1-F)$ for the simulation of a bit-flip channel (see also Appendix C) with fixed parameter $p=0.95$ subjected to noise given by $\hat{\mathcal{B}}$ (see text) with strength $q$ (x-axis), compared to the interleaved building-block strategy to compensate for the noise with perfect and imperfect hardware operation. The yellow line indicates the quality of direct channel simulation subjected to hardware noise, i.e., ^$\hat{\mathcal{E}}^{\prime}_{\text{t}}$. (b) Channel infidelity (Eq. (7)) when transforming an amplitude damping channel of $\gamma=0.1$, Eq. (10), into a depolarizing channel of varying strength (x-axis) by using the interleaved strategy with perfect, and imperfect hardware, the latter modeled in a block fashion (Sec. III.2) with depolarizing noise of strength $q=0.9$.
@@
^$\hat{\mathcal{E}}(\rho)=\sum_{i\in|\textbf{P}_{n}|}p_{i}\,\Sigma_{i}\,\rho\,\Sigma_{i},$ (17)
@@
Any Pauli diagonal channel of the form $\hat{\mathcal{E}}$ can be transformed into a different Pauli diagonal channel by implementing specific post-processing operations. This transformation is achieved by applying imperfect (note these gates are also subjected to Pauli diagonal noise) Pauli gates ^$\hat{\mathcal{E}}(\Sigma_{i}\bullet\Sigma_{i})$ with probability $\lambda_{i}$ after $\hat{\mathcal{E}}$. Mathematically, this transformation can be expressed as
@@
In the particular case where the hardware Pauli diagonal noise corresponds to single qubit depolarizing noise, i.e., ^$\hat{\mathcal{E}}=\hat{\mathcal{D}}$, we observe that $\tilde{p}_{i}=PQ\,\lambda_{i}+(1-QP)/4$, where $P=(4p_{0}-1)/3$ and $Q=(4q_{0}-1)/3$. Consequently, for a target Pauli diagonal channel characterized by $\{\tilde{p}_{i}\}$, given that it satisfies the condition:
@@ VI Conclusions
Our contribution lies in identifying and addressing the key technical ingredients that enable accurate and practical simulations using quantum hardware, including noise tailoring, protocol design, and characterization strategies. Our approach takes a different perspective, rather than mitigating the noise inherent in NISQ devices, we exploit and reshape it to match the noise characteristics of realistic quantum memories, gates, and communication channels. We have introduced tools to tailor local quantum noise, achieving high-fidelity simulations of target channels and demonstrating how this strategy enables the simulation of complete communication protocols on current quantum hardware.
Our method still presents limitations that require further investigation. Apart from the natural limitations of current quantum hardware, the tailoring techniques are currently restricted to single-qubit noise, and while we achieve good accuracy when matching similar noise types, extending to multi-qubit or improving the accuracy for more general noise remains an open challenge. Additionally, although circuit depths remain relatively constant when simulating completely positive maps, measurement overhead—particularly for reconstructing output states or Choi matrices—can be significant. Exploring how to reduce the need for full process tomography –for instance evaluating the Choi matrix directly– is thus an important question for future work.
@@ Appendix A Channel dilatation with an ancillary system
^$\hat{\mathcal{E}}(\rho)=\text{tr}_{\bar{B}}\!\left[\Lambda\left(\left|0\right\rangle\!\left\langle 0\right|_{\bar{A}}\otimes\rho\right)\Lambda^{\dagger}\right],$ (19)
@@
[…]
@@
=(20)${=}\sum_{\begin{subarray}{c}0\leq l,i,j,n<d\\ 0\leq k,m<r\end{subarray}}\rho_{ij}\left\langle l\right|K_{k}\left|i\right\rangle\left\langle j\right|K_{m}^{\dagger}\left|n\right\rangle\left|k,l\right\rangle\!\left\langle m,n\right|$
@@ Appendix B Channel dilatation with an extended qudit
[…]
@@ Appendix C Bit flip channel
^$\hat{\mathcal{E}}(\rho)=P\,\rho+(1-P)X\,\rho\,X.$ (37)
We consider the simulation by implementing the $X$ gate with probability $p$, see Fig. 7a, and by using an ancillary system 7b. We assume that after each gate the system is affected by depolarizing noise, given by

### v2

@@ V Network simulation with noisy quantum devices
In any quantum computing hardware, NISQ devices in particular, noise and decoherence are inevitable factors that affect the reliability of quantum computations. When simulating quantum communication or network processes on such devices, it becomes imperative to incorporate these noisy characteristics. Achieving a faithful simulation of quantum communication processes on quantum hardware necessitates the utilization of diverse tools and techniques. These tools enable the replication of quantum schemes within a quantum circuit while effectively managing -or even leveraging- the inherent noise of the system. As already stressed before, our evaluations typically focus on simulating specific quantum channels, with a designated target denoted as $\hat{\mathcal{E}}_{\text{t}}$, representing the desired channel to be implemented on the quantum hardware. While an ideal, noiseless quantum computer could implement the channel with perfect fidelity using the techniques discussed in Sec. IV, imperfect quantum hardware necessitates the development of strategies to mitigate noise effects, exploit them, or transform certain quantum channels. We primarily rely on static noise models. In practice, however, quantum device noise often exhibits temporal fluctuations due to environmental factors such as temperature drift, vibrations, and electromagnetic interference. While our framework does not simulate noise dynamically at every intermediate time step, it remains applicable by focusing on reproducing the effective initial-to-final quantum process, which can be captured as a single CP map. In this context, we propose several general methods to address these challenges and evaluate their effectiveness through relevant examples.
@@
The first strategy we propose is denoted as the building-block method, illustrated in Fig. 4a. In this method, the circuits used to simulate quantum channels are treated as fixed building blocks. These circuits, representing the fundamental elements of the simulation process, are not directly modifiable, including the sequences of gates and operations. However, it is possible to introduce additional channels within the quantum circuit framework. These supplementary channels, also susceptible to noise, are strategically applied in various configurations. For instance, they can be applied before an input channel, after it, or both (referred to as interleaved), with the aim of achieving a satisfactory approximation of the desired target channel $\hat{\mathcal{E}}_{\text{t}}$}. Note the target channel equals the noiseless implementation of the input channel. Additionally, the method allows for the consideration of non-trivial classical mixtures of these channels applied with different probabilities. It is crucial to note that directly inverting a quantum channel is not possible.
@@
$\hat{\mathcal{E}}_{\text{out}}=\sum_{i,j}p_{ij}\,\hat{\zeta}^{\prime}_{i}\circ\hat{\mathcal{E}}^{\prime}_{\text{t}}\circ\hat{\xi}^{\prime}_{j},$ (16)
such that maximizes the channel fidelity between the output and target channels $F(\hat{\mathcal{E}}_{\text{out}},\hat{\mathcal{E}}_{t})$, such that $\hat{\mathcal{E}}_{\text{out}}\approx\hat{\mathcal{E}}_{\text{t}}$, where $\{p_{ij}\}$ defines a probability distribution and $\{\hat{\zeta}_{i},\hat{\xi}_{j}\}$ are quantum channels. The prime notation, i.e. $\hat{\xi}^{\prime}$ indicates the channel implementation is subjected to computer noise that deviates from the ideal channel implementation $\hat{\xi}$. This deviation can be considered by the method when optimizing over the construction of the channels $\{\hat{\zeta}_{i},\hat{\xi}_{j}\}$. In the examples shown below we model the hardware noise using the models introduced in Sec. III.2.
@@
Instead of solely compensating for noise, this method can also be utilized for channel transformation, with the objective of converting some input channel $\hat{\mathcal{E}}{\text{in}}$ into a –in principle very different– target channel $\hat{\mathcal{E}}{\text{t}}$ using the aforementioned techniques. In the context of large network simulations, there may arise scenarios where a specific type of noise needs to be simulated from a particular point. This strategy enables the direct transformation of the noise present in the simulation into the desired form, effectively harnessing the inherent noise of the quantum computer. From a fundamental perspective, the task of transforming one type of noise into another using additional channels applied before and after holds significant importance, even within the context of noiseless quantum hardware assumptions.
@@
Fig. 5: (a) Channel infidelity (1F)(1-F) for the simulation of a bit-flip channel (see also Appendix C) with fixed parameter $p=0.95$ subjected to noise given by $\hat{\mathcal{B}}$ (see text) with strength $q$ (x-axis), compared to the interleaved building-block strategy to compensate for the noise with perfect and imperfect hardware operation. The yellow line indicates the quality of direct channel simulation subjected to hardware noise, i.e., $\hat{\mathcal{E}}^{\prime}_{\text{t}}$. (b) Channel infidelity (Eq. (7)) when transforming an amplitude damping channel of $\gamma=0.1$, Eq. (10), into a depolarizing channel of varying strength (x-axis) by using the interleaved strategy with perfect, and imperfect hardware, the latter modeled in a block fashion (Sec. III.2) with depolarizing noise of strength $q=0.9$.
@@
$\hat{\mathcal{E}}(\rho)=\sum_{i\in|\textbf{P}_{n}|}p_{i}\,\Sigma_{i}\,\rho\,\Sigma_{i},$ (17)
@@
Any Pauli diagonal channel of the form $\hat{\mathcal{E}}$ can be transformed into a different Pauli diagonal channel by implementing specific post-processing operations. This transformation is achieved by applying imperfect (note these gates are also subjected to Pauli diagonal noise) Pauli gates $\hat{\mathcal{E}}(\Sigma_{i}\bullet\Sigma_{i})$ with probability $\lambda_{i}$ after $\hat{\mathcal{E}}$. Mathematically, this transformation can be expressed as
@@
In the particular case where the hardware Pauli diagonal noise corresponds to single qubit depolarizing noise, i.e., $\hat{\mathcal{E}}=\hat{\mathcal{D}}$, we observe that $\tilde{p}_{i}=PQ\,\lambda_{i}+(1-QP)/4$, where $P=(4p_{0}-1)/3$ and $Q=(4q_{0}-1)/3$. Consequently, for a target Pauli diagonal channel characterized by $\{\tilde{p}_{i}\}$, given that it satisfies the condition:
@@
V.4 Demonstration in real hardware
We demonstrate the practical feasibility of our approach by implementing the basic techniques directly on quantum hardware. This serves as a proof-of-concept for the possibility of tailoring the inherent noise of a quantum computer to engineer specific quantum channels, and eventually whole communication simulations.
Specifically, we realize the circuits shown in Figs. 7 (a) and (b), which correspond to ancilla-assisted simulations of a bit-flip channel and an amplitude damping channel, respectively (see also Secs. C, D). For each case, we compare the naive (theoretically optimal in the absence of noise) parameterization of the rotation angle to a version in which this parameter is variationally optimized in the presence of hardware noise, see Fig. 8. In both cases, we observe that fine-tuning the circuit parameters leads to a better approximation of the target quantum channel.
Experiments were carried out on the IBM Quantum Platform, using the ibm_torino device. For simulation results, we replicated the noise model as faithfully as possible using the real hardware calibration data. The experimental implementations were performed on the same backend for consistency.
Fig. 7: Circuit implementation for real hardware reproduction of (a) Bit-flip channel with strength parameter given by $1-p=\sin^{2}(\theta/2)$, and $R_{y}=\text{exp}\{-i\frac{\theta}{2}Y\}$. (b) Amplitude damping channel with decay parameter $\gamma=\sin^{2}(\theta/2)$. (see also Appendices C, D)
Fig. 8: Demonstration on real hardware using the IBM Quantum Platform ibm_torino device. Realistic backend calibration data is used for the simulation (dashed) curves and real hardware for solid curves. Theoretical best (naive) and variationally tailored (optimized) rotation parameters are considered. Tomographic process reconstruction over 20482048 shots is used in all cases, where channel (Choi) fidelity, compared to the ideal channel implementation, is computed from them. (a) Bit flip channel. (b) Amplitude damping channel.
Our results validate the proposed approach, i.e., even a simple optimization of the rotation parameter results in a significant improvement in channel reproduction fidelity. We note that the overall lower fidelities observed for the amplitude damping in the real hardware case are due to the non-trivial decomposition of the controlled-Y rotation in the machine, which increases circuit depth and cumulative noise. Nevertheless, the optimized case consistently outperforms the naive approach, highlighting the benefit of noise tailoring on real devices.
@@ VI Conclusions
Our contribution lies in identifying and addressing the key technical ingredients that enable accurate and practical simulations using quantum hardware, including noise tailoring, protocol design, and characterization strategies. Our approach takes a different perspective, rather than mitigating the noise inherent in NISQ devices, we exploit and reshape it to match the noise characteristics of realistic quantum memories, gates, and communication channels. We have introduced tools to tailor local quantum noise, achieving high-fidelity simulations of target channels and demonstrating (experimentally on real hardware) how this strategy enables the simulation of complete communication protocols on current quantum hardware.
Our method still presents limitations that require further investigation. Apart from the natural limitations of current quantum hardware, the tailoring techniques are currently restricted to single-qubit noise, and while we achieve good accuracy when matching similar noise types, extending to multi-qubit or improving the accuracy for more general noise, or with additional tools, remains an open challenge. Additionally, although circuit depths remain relatively constant when simulating completely positive maps, measurement overheads can be significant. Exploring how to reduce the need for full process tomography –for instance evaluating the Choi matrix directly– is thus an important question for future work.
@@ Acknowledgments
We acknowledge the use of IBM Quantum Platform services for this work.
@@ Appendix A Channel dilatation with an ancillary system
$\hat{\mathcal{E}}(\rho)=\text{tr}_{\bar{B}}\!\left[\Lambda\left(\left|0\right\rangle\!\left\langle 0\right|_{\bar{A}}\otimes\rho\right)\Lambda^{\dagger}\right],$ (19)
@@
$\Lambda=\begin{pmatrix}\;\boxed{\;\;K_{0}\;\;}\;&\bullet&\bullet&\bullet\;\;\\[5.0pt] \;\boxed{\;\;K_{1}\;\;}\;&\bullet&\bullet&\bullet\;\;\\ \vdots&&&\\[3.0pt] \;\boxed{K_{r-1}}\;&\bullet&\bullet&\bullet\;\;\end{pmatrix}$, (21)
@@
[…]
@@ Appendix B Channel dilatation with an extended qudit
$\Sigma_{i}V_{i}^{\dagger}=\begin{pmatrix}\;\boxed{\;\tilde{K}_{i}\;}\;\;\\[5.0pt] \;\boxed{\;\;\boldsymbol{0}\;\;}\;\;\end{pmatrix}$, (28)
@@ Appendix C Bit flip channel
$\hat{\mathcal{E}}(\rho)=P\,\rho+(1-P)X\,\rho\,X.$ (37)
We consider the simulation by implementing the $X$ gate with probability $p$, see Fig. 9a, and by using an ancillary system 9b. We assume that after each gate the system is affected by depolarizing noise, given by